OpenStudy (anonymous):

Using intercepts to graph linear equations: y = 1/2x - 1 To find y: y = 1/2x - 1 y = 1/2(0) - 1 y = -1 (-1,0) Confused about finding x. If I've done the math correctly then x = -1?

OpenStudy (anonymous):

are you trying to find the x and y-intercepts?

OpenStudy (anonymous):

if so, to find one, set the other variable to zero

OpenStudy (anonymous):

so you did the y-int right

OpenStudy (anonymous):

yes to x and y intercepts "slope and y intercepts"

OpenStudy (anonymous):

so for x, set the y to zero, and solve for the variable x

OpenStudy (anonymous):

when i do that i get x = 1/2 (0, 1/2)

OpenStudy (anonymous):

can you type out your work?

OpenStudy (anonymous):

mine: 0 = (1/2)x -1; add 1 to both sides 1 = (1/2)x; multiply both sides by 2 2 = x

OpenStudy (anonymous):

Thank you, but could you please explain how you got 2. will be working the problem while u type

OpenStudy (anonymous):

and actually, your y-int would be (0,-1)

OpenStudy (anonymous):

your original equation was y = (1/2)x - 1, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you set y to 0, you get the equation 0 = (1/2)x -1; from here on you use algebra to solve for x add 1 to both sides

OpenStudy (anonymous):

to get 1 = (1/2)x multiply both sides by 2 to get 2 = x

OpenStudy (anonymous):

ahh o.k. but what about the -1?

OpenStudy (anonymous):

i added that to both sides in the first step

OpenStudy (anonymous):

so your x-int is (2,0) and your y-int is (0,-1) do you get why?

OpenStudy (anonymous):

OHHHH now i get it: after multiplying y by 0 you get 0 = 1/2x - 1. Next you add one to both sides to get 1 = 1/2x. and when u multiply both sides by 2 you get 2 = x yes i get why now. had to work it in my head and on paper to understand

OpenStudy (anonymous):

technically it's not multiply y by 0, you just do that because you assume that your y coordinate of the x-intercept will be zero, otherwise your point would not fall on the x-axis.

OpenStudy (anonymous):

but everything else is right, im glad you understand

OpenStudy (anonymous):

ty so much for your patience and help.

OpenStudy (anonymous):

you're welcome